40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6136     Entry time: Mon Dec 19 01:54:35 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: another trial of hystersis test 

Another hysteresis test has begun at 1:50 PT, Dec/19.

It will finish after 3 or 4 hours. During the measurement the PSL mechanical shutter will be kept closed.

Time record                       
   Start:  Dec/19 1:50 PST
   End :  Dec/19  5:30 PST
ELOG V3.1.3-